# A Unified Graphical Criterion for Characterizing a Linear Causal Interpretation of Partial Regression Coefficients

### v8

@@ Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies d-separation, single-door, and back-door criteria and proposes a generalized graphical criterion that enables the identification of causal effects not mediated by other explanatory variables. In other words, for almost all parameter points in the parameter space of linear causal models, this criterion is a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results apply to linear models represented by acyclic directed mixed graphs, where the models can include nonlinear terms. The decomposition of linear and nonlinear causal effects is also discussed.
@@ 1 Introduction
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. Among existing approaches, the back-door criterion (Pearl, 1998) and the conditional ignorability assumption (Rubin, 1974) are often regarded as among the most refined conditions. When either of these conditions and some additional assumptions are satisfied, partial regression coefficients in linear structural equation models (SEMs), as developed by Wright (1921), Haavelmo (1943), Jöreskog (1970), Goldberger (1972), and Bollen (1989), can be interpreted as total effects.
However, the following three issues remain unresolved: (1) the control of specific mediated effects; (2) understanding the mechanism underlying post-treatment bias; and (3) causal interpretation in the presence of nonlinear terms. Issues (1) and (2) are fundamental topics in mediation analysis. Since nonlinear terms can be regarded as mediators, a deeper understanding of how conditioning on mediators affects partial regression coefficients is necessary for capturing complex causal relationships.
In relation to issue (1), the single-door criterion (Pearl, 1998) and the sequential ignorability assumption (Imai et al., 2010b) provide partial answers; nevertheless, these conditions cannot accommodate certain mediated effects along specific causal paths. For example, consider the following linear regression equation at the population level:
$Y$ $=\beta_{Y0\mid M_{1}X}+\beta_{YM_{1}\mid X}M_{1}+\beta_{YX\mid M_{1}}X+\epsilon_{Y\mid M_{1}X},$ (1.1)
where $X$, $M_{1}$, and $Y$ are generated by the linear SEM corresponding to Figure 1.
Figure 1: The causal (path) diagram underlying (1.1)
If we are interested in the mediated effect along the path $X\rightarrow M_{2}\rightarrow Y$ in Figure 1, neither condition clearly indicates whether this effect can be identified via the partial regression coefficient $\beta_{YX\mid M_{1}}$. As it turns out, this coefficient still coincides with the causal effect of interest, implying that the existing conditions can be generalized.
@@
As shown in (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce such bias. Other structures that give rise to post-treatment bias are also discussed by Elwert and Winship (2014) and Pearl (2015). Nevertheless, the mechanism underlying post-treatment bias has not yet been fully elucidated, even when restricted to linear models.
In regression analysis, nonlinear terms are often included to capture complex relationships. For example, if the data-generating process of $M_{2}$ is given by $M_{2}=XM_{1}$, the directed edges $X\rightarrow M_{2}$ and $M_{1}\rightarrow M_{2}$ in Figure 1 represent nonlinear causal relationships. In such cases, Imai et al. (2010a) and VanderWeele (2013) examine the identification of total causal effects of $X$ on $Y$ using partial regression coefficients. However, these results rely on specific structural assumptions; therefore, it is necessary to discuss this problem in a more generalized setting.
@@
Against this background, we propose a generalized graphical criterion and provide, in a generic sense, a necessary and sufficient condition under which a partial regression coefficient coincides with a causal effect not mediated by other explanatory variables. In addition, we provide a quantitative characterization of post-treatment bias while controlling for confounding bias. This characterization shows that the joint structure of directed paths and back-door paths is the source of this bias, and that there exist structures that cannot be captured by the conventional concept of path blocking. Finally, we show that the above results can be applied to nonlinear data-generating processes and discuss a decomposition of linear and nonlinear causal effects.
@@
The remainder of this paper is organized as follows. Section 2 introduces key terms as preliminaries. Section 3 formulates the problem addressed in this paper and presents the main theorems. This section also discusses the mechanisms underlying post-treatment bias and applications to nonlinear causal models. All proofs are presented in the Appendix.
@@ 2.1 Model Setup
Assumption 2.1 ensures that the graph associated with model (2.1) is an acyclic directed mixed graph (ADMG), discussed by (Richardson, 2003). With respect to the coefficient matrix $\bm{A}$, strictly speaking, it is sufficient that $\bm{A}$ can be transformed into the structure described in (ii) by reordering the indices of the random variables in $\bm{V}$. However, note that unless the graph structure is known, graphical criteria are in most cases not applicable. In some cases, the graph structure can be identified using methods such as LiNGAM (Shimizu et al., 2006). When the graph structure is unknown, it is necessary to conduct the analysis under each plausible structure and compare the results.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
The discussion in this paper is based on linear models, and the interpretation of bidirected edges is therefore slightly different from that in standard causal diagrams (Pearl, 2009). Accordingly, we distinguish our framework by using the term causal path diagram.
@@
Since the first condition (i) is weaker than that of the back-door criterion, the selective-door criterion allows some directed paths from $X_{j}$ to $X_{i}$ to be blocked. It should be noted, however, that blocking every path from $X_{j}$ to $X_{i}$ except for the directed paths of interest is not sufficient.
Figure 2: An example of unintended post-treatment bias induced by a non-collider
In Figure 2, the set $\{M_{2}\}$ only opens the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
@@ 2.4 Controlled Total Effects
To clarify the interpretation of partial regression coefficients, we define the following partial causal effects.
@@
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}=\bm{z}),\,do(X_{j}=x_{j})],$ (2.4)
@@ 3.1 Problem Formulation
Suppose we are interested in the causal effect of $X_{j}$ on $X_{i}$. Let $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$ denote the set of explanatory variables, and consider the following regression equation at the population level:
@@
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$ denotes a set of covariates. In this paper, the partial regression coefficient is defined using ordinary covariances, not conditional covariances. See the Appendix for details. The objective is to derive a sufficient condition on $\bm{Z}$ under which $\beta_{ij\mid\bm{Z}}=\tau_{ij\mid do(\bm{Z})}$ holds. If the coefficient $\beta_{ij\mid\bm{Z}}$ corresponds to the causal effect, then, under standard regularity conditions for random explanatory variables, the least squares estimator provides a consistent estimate of this effect.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on post-treatment variables (i.e., descendants in the causal diagram) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
@@
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ opens a back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
$\bm{S}_{2}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ blocks every back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking every back-door path. The first condition (i) of the selective-door criterion implies that $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. In particular, conditioning on a descendant $X_{p}$ of the outcome variable $X_{i}$ is problematic, since whenever there exists an opened directed path $X_{j}\rightarrow\rightarrow X_{i}$, a sequence of vertices $X_{j}\rightarrow\rightarrow X_{i}\rightarrow\rightarrow X_{p}\leftarrow\leftarrow X_{i}$ exists, leading to post-treatment bias. On the other hand, if each path from $X_{j}$ to $X_{i}$ is blocked (i.e., $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ ), the partial regression coefficient vanishes even when a decendant of the outcome variable $X_{i}$ is included among the conditioning variables.
@@ 3.4 Necessity for Almost All Parameter Points
3.5 Application to Nonlinear Data-Generating Processes
Since simple linear models have limitations in their approximation performance, we consider the following structural equation model:
$X_{i}=\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\sum_{k}a_{h_{ik}}h_{ik}(\bm{PA}(X_{i}))+u_{i},$ (3.3)
where each $h_{ik}$ is a known nonlinear function of $\bm{PA}(X_{i})$, and each term $h_{ik}(\bm{PA}(X_{i}))$ can capture interactions and effect modification (see VanderWeele (2009)). This model accommodates spline-based models (Cox, 1972; de Boor, 1972; Wahba, 1978; Silverman, 1985; Eilers and Marx, 1996). Even when nonlinear terms are present, the discussion so far remains applicable by explicitly representing them as vertices in the graph and replacing the directed edges corresponding to nonlinear causal relationships with bidirected edges. In the following, for the sake of clarity, we illustrate this point through a concrete example.
To accommodate nonlinear structures, we impose an additional standard assumption on the causal path diagram.
Assumption 3.8.
Any random vector consisting only of error terms corresponding to vertices that are not adjacent to $X_{i}$ by a bidirected edge is independent of $u_{i}$.
$\begin{cases}Z=u_{Z}\\ X=a_{XZ}Z+u_{X}\\ M=a_{MX}X+a_{MZ}Z+u_{M}\\ H=h(X,M)=u_{H}\\ Y=a_{YX}X+a_{YM}M+a_{YH}H+a_{YZ}Z+u_{Y},\end{cases}$ (3.4)
Figure 3: The causal path diagram associated with (3.4)
Assumption 3.8 guarantees the absence of correlation between $u_{H}$ and $u_{Y}$ when $u_{Z}$, $u_{X}$, and $u_{M}$ are independent of $u_{Y}$. Depending on the nonlinear function $h$, some of the bidirected edges in Figure 3 may be omitted, since uncorrelatedness may hold. This causal path diagram shows that a bidirected edge does not necessarily indicate the presence of a confounding factor; it may also represent a causal relationship that cannot be expressed by path coefficients. For instance, the paths $X\leftrightarrow H\to Y$ and $X\to M\leftrightarrow H\to Y$ mediate the nonlinear causal effect of $X$ on $Y$. Since model (3.3) can be described as a special case of model (2.1), as in model (3.4), all the results obtained thus far are applicable to the identification of linear causal effects.
In this case, the following decomposition of the “true” total effect holds:
$\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]$ $=\tau_{YX}+a_{YH}\frac{d}{dx}\,\mathrm{E}[h(X,M)\mid do(X=x)],$
where $\tau_{YX}=a_{YX}+a_{YM}a_{MX}$ is the total effect in the sense of linear SEM. More generally, this decomposition is expressed as
$\tau_{YX}\coloneq\frac{\partial}{\partial x}\,\mathrm{E}[Y\mid do(\bm{H}=\bm{h}),\,do(X=x)],$
$\delta_{YX}(x)\coloneq\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]-\tau_{YX},$
$\frac{d}{dx}\,\mathrm{E}[Y\mid do(X=x)]=\tau_{YX}+\delta_{YX}(x),$
where $\bm{H}$ denotes all nonlinear terms. The second term $\delta_{YX}(x)$ represents the causal effects originated from nonlinear structures and includes the mediated interactive effect (VanderWeele, 2013). If the nonlinear terms are not differentiable, one can instead consider the difference between two intervention levels, $do(X=x)$ and $do(X=x^{\prime})$. In general, the identifiability of the nonlinear causal effect $\delta_{YX}(x)$ depends on the form of nonlinear functions and causal relationships among the ancestors of $Y$.
Here, we restrict our discussion to the case of (3.4). Note that the derivative constituting the nonlinear effect can be expressed as follows.
$\frac{d}{dx}\,\mathrm{E}[h(X,M)\mid do(X=x)]=\frac{d}{dx}\,\mathrm{E}\left[\tilde{h}\left(x,a_{MX},a_{MZ}Z+u_{M}\right)\right],$ (3.5)
where $\tilde{h}$ is a function. Fortunately, this nonlinear causal effect of $X$ on $H$ can be identified because $a_{MX}=\beta_{MX\mid Z}$ and $a_{MZ}Z+u_{M}=M-\beta_{MX\mid Z}X$ hold in the following regression equation:
$M=\beta_{M0\mid XZ}+\beta_{MX\mid Z}X+\beta_{MZ\mid X}Z+\epsilon_{M\mid XZ}.$
Moreover, $\tau_{YX}=\beta_{YX\mid HZ}$ and $a_{YH}=\beta_{YH\mid XZ}$ hold in the following regression equation:
$Y=\beta_{Y0\mid XHZ}+\beta_{YX\mid HZ}X+\beta_{YH\mid XZ}H+\beta_{YZ\mid XH}Z+\epsilon_{Y\mid XHZ}.$
Therefore, the effects $\tau_{YX}$, $\delta_{YX}(x)$ and $\tau_{YX}+\delta_{YX}(x)$ are identifiable if we can also observe $M$ and $Z$.
@@ 4 Discussion
We have derived generalized theorems that encompass existing results characterizing partial regression coefficients based on conventional graphical criteria. In causal diagrams represented by ADMGs, bidirected edges indicate the presence of unobserved common causes. In contrast, in the causal path diagrams defined in this paper, bidirected edges represent only correlations among the error terms, which constitutes a major difference. As in the case of nonlinear terms, for graphs with cycles or for partial ancestral graphs (Zhang, 2008), the results of this paper remain applicable as long as the graphs can be transformed into an ADMG while disregarding the usual interpretation of bidirected edges. This is largely due to the fact that the all proofs of the theorems in this paper rely only on algebraic arguments.
In the standard linear SEM, where bidirected edges represent the presence of unobserved confounders, the linear causal interpretation of partial regression coefficients under Theorem 3.1 can be summarized as follows:
$\beta_{ij\mid\bm{Z}}$ $=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,do(\bm{Z}^{j}=\bm{z}^{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,\bm{Z}^{j}=\bm{z}^{j}]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j})],$
where $\bm{Z}_{j}\coloneq\bm{Z}\cap\bm{DE}(X_{j})$, and $\bm{Z}^{j}\coloneq\bm{Z}\setminus\bm{Z}_{j}$. By definition, it is clear that $\tau_{ij\mid do(\bm{Z})}\equiv\tau_{ij\mid do(\bm{Z}_{j})}$. In particular, when $\bm{Z}_{j}\neq\emptyset$, special attention must be paid to the following relationship.
$\beta_{ij\mid\bm{Z}}$ $\not\equiv\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),\,\bm{Z}=\bm{z}].$
Here, $\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})$ denotes intervening on $X_{j}$ after observing $\bm{Z}=\bm{z}$, and is conceptually similar to the counterfactual query $Q_{c}$ (Pearl, 2015). If the back-door criterion is satisfied, the symbol $\not\equiv$ can be replaced with ==. These interpretations rely on the absence of interaction or effect modification, and therefore does not necessarily hold when nonlinear terms are present. See the Appendix for details.
As a direction for future research, a more detailed investigation of the estimation of decomposed nonlinear causal effects is required. In particular, the discussion of identifiability of nonlinear causal effects in this paper is limited to whether the causal parameter (or its function) can be expressed as a function of the observed distribution. For example, the interchangeability of differentiation and expectation in (3.5) is expected to affect estimation accuracy. Therefore, further analysis under appropriate assumptions on the nonlinear terms remains an important topic for future work.

### v9

@@ Abstract
This paper characterizes the values of partial regression coefficients—defined as projection coefficients onto the space spanned by explanatory variables—for random variables generated by linear structural equation models using graphical structures. First, we derive a generalized graphical criterion that unifies the d-separation, single-door, and back-door criteria. This criterion provides a generically necessary and sufficient condition under which a partial regression coefficient coincides with the linear causal effect not mediated by other explanatory variables. Second, we reveal the mechanism underlying post-treatment bias and characterize it quantitatively. This provides a unified framework for discussing the graph structures that generate post-treatment bias, which have previously been examined individually, and clarifies the existence of graph structures that cannot be prevented by the conventional concept of path-blocking. These results are based on the algebraic properties of acyclic directed mixed graphs and do not rely on any specific probability distribution, making them applicable to a broad class of linear models.
@@ 1 Introduction
The conditions under which partial regression coefficients, defined as projection coefficients onto the space spanned by explanatory variables, admit a causal interpretation have been discussed in various studies. For example, the conditional ignorability (independence) assumption (Rubin, 1974; Angrist and Pischke, 2009) and the back-door criterion (Pearl, 1998) are often regarded as the most refined conditions. When either of these conditions and some additional assumptions are satisfied, partial regression coefficients in linear structural equation models (SEMs), as developed by Wright (1921), Haavelmo (1943), Jöreskog (1970), Goldberger (1972), and Bollen (1989), can be interpreted as total effects.
However, these conditions presuppose that the quantities of interest are total effects and therefore exclude mediators from the explanatory variables. Consequently, in situations such as that shown in Figure 1, where confounding bias can be eliminated only by conditioning on a mediator, we cannot causally interpret the partial regression coefficients based on these conditions.
Figure 1: An example where the back-door criterion is not applicable
In such cases, it is known that the d-separation and single-door criteria (Spirtes et al., 1998; Pearl, 1998) and their simple extensions are applicable to identify the effect that is not mediated by other explanatory variables. To illustrate this point, consider the following linear regression model at the population level:
$Y$ $=\beta_{Y0\mid M_{1}X}+\beta_{YM_{1}\mid X}M_{1}+\beta_{YX\mid M_{1}}X+\epsilon_{Y\mid M_{1}X},$
where the coefficients are defined so as to minimize the expected squared value of the residual $\epsilon_{Y\mid M_{1}X}$. In this regression equation, the coefficient $\beta_{YX\mid M_{1}}$ coincides with the effect not mediated by $M_{1}$, that is, the effect transmitted along the path $X\rightarrow M_{2}\rightarrow Y$. Thus, even when the total effect is not identifiable, it may still be possible to identify effects transmitted along some paths, which can provide further insight into the total effect. Therefore, understanding specific components of the causal structure is important not only in mediation analysis, where particular mediation effects are of interest.
@@
As illustrated by the regression equation above, it is sometimes possible to identify an effect that excludes a particular mediation effect by appropriately including mediators among the explanatory variables. On the other hand, it is well known that conditioning on a mediator may induce unintended post-treatment bias. For example, a variable such as $M_{2}$ in Figure 1, which is affected by multiple variables, is referred to as a collider, and it is known that including such a variable among the explanatory variables can induce this type of bias. Other cases that give rise to post-treatment bias have been discussed in, for example, Elwert and Winship (2014), Pearl (2015), and Cinelli et al. (2024). However, the mechanisms underlying post-treatment bias have not been fully elucidated.
@@
Against this background, we propose a generalized graphical criterion that unifies the d-separation, single-door, and back-door criteria. Based on the proposed criterion, we provide a generically necessary and sufficient condition under which a partial regression coefficient coincides with a causal effect not mediated by other explanatory variables. This criterion provides an important insight: it shows that the analogy to the single-door criterion—namely, that the causal effect of interest can be identified by blocking all paths except the target causal path—does not always hold. In addition, we provide a quantitative characterization of post-treatment bias while controlling for confounding bias, which captures the mechanisms underlying the post-treatment biases discussed in previous studies. This characterization reveals the existence of structures that cannot be explained by the conventional concept of path blocking and clarifies why the first condition of the criterion we propose is required.
To clarify the contribution of this study, we review previous work on the identification of linear SEMs using graphical approaches. The single-door and back-door criteria (Spirtes et al., 1998; Pearl, 1998) enable the identification of direct and total effects, respectively, while the front-door criterion (Pearl, 1995, 2009) is applicable to the identification of total effects through the product of partial regression coefficients. Extensions of instrumental variable approaches were proposed by Brito and Pearl (2002a), Brito and Pearl (2002b), and Chan and Kuroki (2010), while methods alternative to the instrumental variable approach were studied by Stanghellini and Pakpahan (2015). Tian (2009) proposed a systematic procedure for identifying individual direct effects in a class of linear SEMs by reducing the identification problem to a maximum flow problem on a graphical structure. Foygel et al. (2012) proposed the half-trek criterion, which provides a flexible framework for identifying model parameters. This criterion was further extended by Chen (2015), Chen et al. (2016) and Barber et al. (2022).
Despite these advances, existing studies primarily focus on the identification of direct effects, total effects, or the entire set of model parameters. In contrast, the identification of effects mediated only through a subset of paths has received relatively little attention. This paper not only strengthens the causal interpretation of partial regression coefficients but also serves to bridge this gap.
@@
The remainder of this paper is organized as follows. Section 2 introduces key terms as preliminaries. Section 3 formulates the problem addressed in this paper and presents the main theorems. This section also discusses the mechanisms underlying post-treatment bias. All proofs are presented in the Appendix.
@@ 2.1 Model Setup
Assumption 2.1 ensures that the graph associated with model (2.1) is an acyclic directed mixed graph (ADMG), discussed by (Richardson, 2003). With respect to the coefficient matrix $\bm{A}$, strictly speaking, it is sufficient that $\bm{A}$ can be transformed into the structure described in (ii) by reordering the indices of the random variables in $\bm{V}$. However, note that unless the graph structure is known, graphical criteria are in most cases not applicable. When the graph structure is unknown, it is necessary to conduct the analysis under each plausible structure and compare the results.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
We define the graph under consideration as follows.
@@
Since the first condition is weaker than that of the back-door criterion, the selective-door criterion allows some directed paths from $X_{j}$ to $X_{i}$ to be blocked. It should be noted, however, that blocking every path from $X_{j}$ to $X_{i}$ except for the directed paths of interest is not sufficient.
Figure 2: An example where a non-collider induces unintended post-treatment bias
In Figure 2, the set $\{M_{2}\}$ only opens the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking. We will examine in detail in a later section how this structure induces unintended post-treatment bias.
@@ 2.4 Controlled Total Effects
To clarify the interpretation of partial regression coefficients, we define the following causal effect.
@@
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}=\bm{z}),\,do(X_{j}=x_{j})],$
@@ 3.1 Problem Formulation
Suppose we are interested in the causal effect of $X_{j}$ on $X_{i}$. Let $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$ denote the set of explanatory variables, and consider the following linear regression equation at the population level:
@@
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$ denotes a set of covariates, and the coefficients are projection coefficients onto the space spanned by explanatory variables. See the Appendix for details. The objective is to derive a sufficient condition on $\bm{Z}$ under which $\beta_{ij\mid\bm{Z}}=\tau_{ij\mid do(\bm{Z})}$ holds. If the coefficient $\beta_{ij\mid\bm{Z}}$ corresponds to the causal effect, then, under standard regularity conditions for random explanatory variables, the least squares estimator provides a consistent estimate of this effect.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on post-treatment variables (i.e., descendants in the graph) are collectively referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why the first condition of the selective-door criterion is required.
@@
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ opens a back-door path $X_{k}\dashleftarrow X_{i}$}.\,\},$
$\bm{S}_{2}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ blocks every back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}.$
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking every back-door path. The first condition of the selective-door criterion implies that $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias.
Figure 3 illustrates four well-known structures that induce post-treatment bias. In (a) and (b), as considered under path blocking, including only the collider $Z$ as a covariate leads to bias. In contrast, in (c) and (d), $Z$ is not regarded as a collider, yet it is known to generate post-treatment bias; Cinelli et al. (2024) refers to such a variable as a virtual collider. Although these structures are often discussed individually, as Theorem 3.6 shows, they can all be uniformly described as a combined structure of opened directed paths and opened backdoor paths.
The reason the single-door criterion excludes descendants of the outcome variable from the set of covariates is that it accounts for the possibility of structures such as (c). Note that this criterion assumes the blocking of all directed paths except for the directed edge; hence, structures such as (d) need not be considered. Furthermore, under the d-separation criterion, all directed paths are blocked, implying that including descendants of the outcome variable as covariates does not cause bias.
Figure 3: Graph structures inducing post-treatment bias
@@ 4 Discussion
We have derived generalized theorems that encompass existing results characterizing partial regression coefficients based on conventional graphical criteria. In this paper, the bidirected edges in an ADMG do not necessarily represent the presence of a confounding variable with a linear causal relationship; rather, they allow for arbitrary dependency structures. Therefore, as long as a graph can be transformed into an ADMG, the results presented here can be directly applied to other classes of graphs.
The characterization of post-treatment bias in this paper assumes that confounding bias has been eliminated, so deriving a general expression that also incorporates confounding bias remains a task for future research. Obtaining such an expression is expected to facilitate the discovery of new estimators for causal parameters expressed as functions of partial regression coefficients.